A 10-π‘˜π‘” box starting from rest is pulled using a rope that makes an angle of 30o with the horizontal. If it travels a distance of 10 π‘˜π‘š in 2𝑠, what is the magnitude of the force exerted by the rope?


1
Expert's answer
2021-10-03T13:25:44-0400

Write the equation that ties together all forces by Newton's second law:


"ma=F\cos\theta,\\\space\\
F=\frac{ma}{\cos\theta}."

The acceleration can be expressed as

"a=\frac{2d}{t^2}.\\\space\\
F=\frac{2md}{t^2\cos\theta}=58\text { kN }"